A uniform horizontal beam, 6m long is supported at points A and B, 1m and 4m respectively, from one end, If the supporting force at B is 400N, calculate the mass of the beam and the supporting force at A.
1
Expert's answer
2020-11-16T07:51:54-0500

For the equilibrium:

"W=Mg=R_a+R_b"

Taking moments about point A:

"W(0.5(6)-1)=R_b(4-1)\\W=1.5(400)=600\ N"

"m=\frac{W}{g}=\frac{600}{9.8}=61\ kg"

"R_a=600-400=200\ N"